OpenStudy (anonymous):

solve the interval (0,2 pi) 10cox^2x +3sinx-9=0

OpenStudy (anonymous):

Use the identity: cos^2 x = 1 - sin^2 x 10(1 - sin^2 x) + 3sin x - 9 = 0 10(sin^2 x) - 3 sinx - 1 = 0 sin x = 1/2 or -1/5

OpenStudy (anonymous):

So, in degrees, x = 30, 150, 348.46, 191.54